OpenStudy (anonymous):

how do you Divide (y-4)(y+3)/(y-2)(y-4) by (y+5)(y-3)/(y-2)(y-5)

OpenStudy (turingtest):

the rule is\[\huge{\frac ab\over\frac cd}=\frac ab\cdot\frac dc\]so flip the bottom fraction, then just multiply across (before you distribute note that some things will cancel)

OpenStudy (anonymous):

by she means times or division again?